# Inference for Deep Neural Network Estimators in Generalized Nonparametric Models

### v3

@@ A.1 Binomial Regression
$\mathbb{P}(y_{i}=k|\mathbf{x}_{i})={\binom{n_{\text{trial}}}{k}}p^{k}\big(\mathbf{x}^{(i)}\big)\big(1-p\big(\mathbf{x}^{(i)}\big)\big)^{k},$
@@ A.4 Additional Experiments on Dimensionality and Nonlinearity
As shown in Table S.7, increasing the dimension to $p=20$ degrades estimation accuracy, with $\text{MAE}_{f}$ rising from 0.28 to 0.55 and empirical standard deviation from 0.34 to 0.62 under the logistic regression model with $n=700$. This reflects increased bias and variability, leading to a drop in coverage from 93.7% to 92.1%. Similar effects of high-dimensional noise on random forests were observed in Wager & Athey (2018). Increasing the sample size to $n=2000$ under $p=20$ improves estimation, restoring coverage to nearly 95%. Conversely, reducing the dimension to $p=7$ improves neural network performance (Table S.8), with $\text{MAE}_{f}$ decreasing to 0.23 and empirical standard deviation to 0.27. The reduced bias yields shorter intervals and coverage close to 95%.
@@ A.6 Additional Experiments on $r$
We also evaluate computational cost across different $r$ values by recording the training time of a single neural network. All the simulations are conducted in the institute high performance computing platform. The simulation is repeated 300 times, with $B$ networks trained per repetition. As shown in Table S.12, training one network takes only a few seconds, while the main cost arises from repeatedly training the $B$ networks. Following Wager & Athey (2018) and Fei & Li (2024), we let $B$ grow proportionally with $n$ and apply bias correction to maintain variance estimation accuracy, largely reducing computation. Although $300\times B$ trainings would otherwise be intensive, high-performance computing with parallel nodes and multi-CPU systems makes the experiments feasible.
@@ B.3 Additional Experiments on Cross-Site Model Transferability
$y_{i}$ Response variable $\mathrm{Var}[\cdot]$ Variance operator
$\mathbf{z}_{i}$ $(\mathbf{x}_{i},y_{i})$ $V_{k}$ $\mathrm{Var}(T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}))$
@@
$T_{k}(\mathbf{z}_{1},...,\mathbf{z}_{k})$ Hoeffding decomposition in (C.5) $\widehat{\mathrm{Var}}_{i}$ $\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}/(B-1)$
@@ C.1 Preliminaries
C.3 Proof of Corollary 3.6
According to Theorem 3.5, we have that when $\Delta_{n}(\widehat{f}^{b_{j}})\leq C\phi_{r}L\log^{2}r$,
$R_{n}(\widehat{f}^{b_{j}},f_{0})\precsim\phi_{r}L\log^{2}r.$
Lemma C.4 further indicates that
$R_{n}(\widehat{f}^{b_{j}},f_{0})\geq\frac{c}{2}\mathbb{E}(\widehat{f}^{b_{j}}(\bm{X})-f_{0}(\bm{X}))^{2}.$
Here, $\bm{X}$ comes from the distribution $\mathbb{P}_{\mathbf{x}}$ independently, and the expectation is with respect to $\bm{X}$. Hence we have
$\mathbb{E}(\widehat{f}^{b_{j}}(\bm{X})-f_{0}(\bm{X}))^{2}\leq C\phi_{r}L\log^{2}r$
for some constant $C>0$. By the inequality of $\mathbb{E}(\mathbb{E}(X|U)-U)^{2}\leq\mathbb{E}(X-U)^{2}$ for any random variables $X$ and $U$, we have the expectation
$\mathbb{E}(\mathbb{E}\widehat{f}^{b_{j}}(\bm{X})|\bm{X}-f_{0}(\bm{X}))^{2}\leq C\phi_{r}L\log^{2}r\to 0$
for some constant $C>0$. By the assumption $\lim_{n\to\infty}n\phi_{r}L\log^{2}(r)/(r^{2}\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x}))=0$ in Assumption 3.4, we have
$\frac{n}{r^{2}\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x})}\cdot\mathbb{E}\big|\mathbb{E}(\widehat{f}^{B}(\bm{X})-f_{0}(\bm{X})|\bm{X})\big|^{2}~\to~0.$
This completes the proof of Corollary 3.6.
@@
By Assumption 3.4, recall $\gamma>\frac{1}{1+\min_{j=0,...,q}(2\gamma_{j}^{*})/(2\gamma_{j}^{*}+t_{j})}$. For the sufficiently small $\varepsilon>0$, it is easy to see that
@@
Let $W_{b_{j}}$ be the number of subsamples $\mathcal{I}^{b_{j}}$ that contain index $i$, and define $g_{1,r}(\mathbf{z};\mathbf{x}_{*})=\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z},\mathbf{z}_{2},...,\mathbf{z}_{r})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. We have $W_{b_{j}}$ is independent of $\mathbf{z}_{i}$ and $\mathbb{E}W_{b_{j}}=B{\binom{n-1}{r-1}}/{\binom{n}{r}}$. Hence, the equation above involving the term $\mathbf{z}_{i}$ can be further expressed as
@@
$\frac{\sqrt{n}}{r\cdot\sqrt{\mathrm{Var}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))}}\cdot\mathring{f}^{B}(\mathbf{x}_{*})~{\stackrel{{\scriptstyle d}}{{\rightarrow}}}~\mathcal{N}(0,1).$
By the definition of $\xi_{1,r}(\mathbf{x}_{*})$, it is easy to see that $\mathrm{Var}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))=\xi_{1,r}(\mathbf{x}_{*})$, hence we have
@@
$\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big)$ $=\frac{1}{{\binom{n}{r}}}\Bigg({\binom{n-1}{r-1}}\sum_{i=1}^{n}T_{1}(\mathbf{z}_{i})+{\binom{n-2}{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})$
@@
$=\frac{1}{{\binom{n}{r}}}\Bigg({\binom{n-2}{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})+...+\sum_{i_{1}<...<i_{r}}T_{r}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{r}})\Bigg).$
Denote by $V_{k}=\mathrm{Var}(T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}))$, we then calculate the value of $\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$, which we will show later this term is negligible. To calculate this term, we apply the equality $\mathbb{E}X_{n}^{2}=\mathbb{E}\mathrm{Var}(X_{n}|Y_{n})+\mathbb{E}(\mathbb{E}X_{n}|Y_{n})^{2}$. Then the term $\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ can be written as
$\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ $=\underbrace{\mathbb{E}\big(\mathrm{Var}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big)\big)}_{J_{1}}$
@@
$J_{2}$ $=\frac{1}{{\binom{n}{r}}^{2}}\cdot\sum_{k=2}^{r}{\binom{n-k}{r-k}}^{2}\cdot{\binom{n}{k}}V_{k}$
$=\sum_{k=2}^{r}\frac{(n-k)!r!}{n!(r-k)!}{\binom{r}{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\cdot\sum_{k=2}^{r}{\binom{r}{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})).$
@@
$J_{1}$ $=\frac{1}{B^{2}}\sum_{k=2}^{r}\mathrm{Var}(W_{i_{1}...i_{k}}){\binom{n}{k}}V_{k}$
$\leq\frac{1}{B}\sum_{k=2}^{r}\frac{{\binom{n-k}{r-k}}}{{\binom{n}{r}}}{\binom{n}{k}}V_{k}$
$=\frac{1}{B}\sum_{k=2}^{r}{\binom{r}{k}}V_{k}$
$\leq\frac{\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))}{B}.$
Here, the first inequality is by $\mathrm{Var}(W_{i_{1}...i_{k}})\leq B\cdot\frac{{\binom{n-k}{r-k}}}{{\binom{n}{r}}}$ and the uncorrelation on different terms $T_{k}$, and the last inequality is by $\sum_{k=2}^{r}{\binom{r}{k}}V_{k}\leq\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$. We hence conclude that
@@
$\leq\bigg(\frac{r(r-1)}{n(n-1)}+\frac{1}{B}\bigg)\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})),$
@@
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ $\leq\bigg(\frac{1}{n\xi_{1,r}(\mathbf{x}_{*})}+\frac{n}{Br^{2}\xi_{1,r}(\mathbf{x}_{*})}\bigg)\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$
@@
Here, we utilize the fact that $\widehat{f}^{b}(\mathbf{x}_{*})$ is bounded. The bound of $\widehat{f}^{b}(\mathbf{x}_{*})$ indicates that $\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},\ldots,\mathbf{z}_{r}))$ is bounded. Additionally, we rely on Assumption 3.4 that $\liminf_{n\to+\infty}n\xi_{1,r}(\mathbf{x}_{*})\geq n^{1-\varepsilon-\gamma}\to+\infty$ and $\liminf_{n\to+\infty}Br^{2}\xi_{1,r}(\mathbf{x}_{*})/n\to+\infty$. From (C.6), we have
@@
Without loss of generality, we calculate the term $\widehat{V}_{1}$, and the terms $\widehat{V}_{i}$ with $i\neq 1$ are exactly the same. Consider $\widehat{V}_{1}^{2}$ and $B/n\to c$ for some constant $c>0$. Then by the inspiration of the proof of Lemma 13 of Wager & Athey (2018), we can define
@@
$\quad+\frac{r}{n}\sum_{k=1}^{r-1}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k+1}}}{{\binom{n}{k+1}}}\bigg)\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k+1}(\mathbf{z}_{1},\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}})$
$\quad+\frac{r}{n}\sum_{k=2}^{r}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k}}}{{\binom{n}{k}}}\bigg)\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}),$
@@
Further denote by $\widehat{\mathrm{Var}}_{i}=\frac{1}{B-1}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}$, we have
$\widehat{R}_{B}=\frac{n(n-1)}{(n-r)^{2}}\frac{1}{B}\sum_{i=1}^{n}\widehat{\mathrm{Var}}_{i}.$
@@
$R_{B}-\widehat{R}_{B}=\frac{n(n-1)}{(n-r)^{2}}\bigg(\underbrace{\sum_{i=1}^{n}2M_{i}\overline{\varepsilon}_{i}}_{\Delta_{1}}+\underbrace{\sum_{i=1}^{n}\overline{\varepsilon}_{i}^{2}-B^{-1}\mathrm{Var}_{i}}_{\Delta_{2}}-\underbrace{\frac{1}{B}\sum_{i=1}^{n}\widehat{\mathrm{Var}}_{i}-\mathrm{Var}_{i}}_{\Delta_{3}}\bigg).$
Here, $\mathrm{Var}_{i}=\mathrm{Var}_{b}(Z_{b_{j}i})$ with variance condition on the dataset $\mathcal{D}_{n}$. We bound the three terms one by one. For the term $\Delta_{1}$, by the definition of $M_{i}$ and $\overline{\varepsilon_{i}}$, we have
@@
Conditional on the dataset $\mathcal{D}_{n}$, $\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}$ is independent with different $b_{j}$. It is easy to see that $\mathbb{E}_{b}\Delta_{1}=0$, and we give the calculation of $\mathrm{Var}_{b}\Delta_{1}$. From the expression of $\Delta_{1}$, conditional on data set $\mathcal{D}_{n}$, $\varepsilon_{b_{j}i}$ are independent for different $b_{j}$, it is easy to see that
$\mathrm{Var}_{b}\Delta_{1}\asymp\frac{\mathrm{Var}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}}{B}.$
@@
$\mathrm{Var}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}=\sum_{i=1}^{n}M_{i}^{2}\mathrm{Var}_{b}\varepsilon_{b_{j}i}+\sum_{i\neq i^{\prime}}M_{i}M_{j}\operatorname{\rm Cov}_{b}(\varepsilon_{b_{j}i},\varepsilon_{b_{j}i^{\prime}}).$
From the definition of $M_{i}$, we can easily bound $M_{i}=O(r/n)$, for the covariance of $\varepsilon_{b_{j}i}$ and $\varepsilon_{b_{j}i^{\prime}}$, simple calculation of the subsampling randomness gives us that $\mathrm{Var}_{b}\varepsilon_{b_{j}i}=O(\frac{r}{n}(1-\frac{r}{n})=O(r/n)$, and $\operatorname{\rm Cov}_{b}(\varepsilon_{b_{j}i},\varepsilon_{b_{j}i^{\prime}})=O(|\frac{r(r-1)}{n(n-1)}-\frac{r^{2}}{n^{2}})|=O(\frac{r}{n^{2}})$, we hence have
$\mathrm{Var}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}$ $=O\bigg(n\cdot\frac{r^{2}}{n^{2}}\cdot\frac{r}{n}+n^{2}\cdot\frac{r^{2}}{n^{2}}\cdot\frac{r}{n^{2}}\bigg)$
@@
This indicates that $\mathrm{Var}_{b}\Delta_{1}=O(\frac{r^{3}}{Bn^{2}})$. Combining the fact that $\mathbb{E}_{b}\Delta_{1}=0$, we conclude that $|\Delta_{1}|=O_{p}(\frac{r^{3/2}}{n\sqrt{B}})$.

### v4

@@ A.1 Binomial Regression
$\mathbb{P}(y_{i}=k|\mathbf{x}_{i})={n_{\text{trial}}\choose{k}}p^{k}\big(\mathbf{x}^{(i)}\big)\big(1-p\big(\mathbf{x}^{(i)}\big)\big)^{k},$
@@ A.4 Additional Experiments on Dimensionality and Nonlinearity
As shown in Table S.7, increasing the dimension to $p=20$ degrades estimation accuracy, with $\text{MAE}_{f}$ rising from 0.28 to 0.55 and empirical standard deviation from 0.34 to 0.62 under the logistic regression model with $n=700$. This reflects increased bias and variability, leading to a drop in coverage from 93.7% to 92.1%. Similar effects of high-dimensional noise on random forests were observed in Wager and Athey (2018). Increasing the sample size to $n=2000$ under $p=20$ improves estimation, restoring coverage to nearly 95%. Conversely, reducing the dimension to $p=7$ improves neural network performance (Table S.8), with $\text{MAE}_{f}$ decreasing to 0.23 and empirical standard deviation to 0.27. The reduced bias yields shorter intervals and coverage close to 95%.
@@ A.6 Additional Experiments on $r$
We also evaluate computational cost across different $r$ values by recording the training time of a single neural network. All the simulations are conducted in the institute high performance computing platform. The simulation is repeated 300 times, with $B$ networks trained per repetition. As shown in Table S.12, training one network takes only a few seconds, while the main cost arises from repeatedly training the $B$ networks. Following Wager and Athey (2018) and Fei and Li (2024), we let $B$ grow proportionally with $n$ and apply bias correction to maintain variance estimation accuracy, largely reducing computation. Although $300\times B$ trainings would otherwise be intensive, high-performance computing with parallel nodes and multi-CPU systems makes the experiments feasible.
@@ B.3 Additional Experiments on Cross-Site Model Transferability
$y_{i}$ Response variable $\operatorname{{\rm Var}}[\cdot]$ Variance operator
$\mathbf{z}_{i}$ $(\mathbf{x}_{i},y_{i})$ $V_{k}$ $\operatorname{{\rm Var}}(T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}))$
@@
$T_{k}(\mathbf{z}_{1},...,\mathbf{z}_{k})$ Hoeffding decomposition in (C.5) $\widehat{\operatorname{{\rm Var}}}_{i}$ $\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}/(B-1)$
@@ C.1 Preliminaries
C.3 Proof of Proposition 3.6
We adapt the proof technique in Schmidt-Hieber (2020) to construct the approximation error in such class of ReLU networks. Suppose that
$\phi_{r}=r^{-\frac{2m}{2m+t}},$
and let $c=\frac{2m}{2m+t}$. Then we specifically construct the neural network $\widetilde{f}$ with the approximation rate
$\sup_{f_{0}\in\mathcal{G}}\|\widetilde{f}-f_{0}\|_{\infty}\precsim\phi_{r}^{\frac{1+\delta(t+2m)/t}{2}}.$
As we change the sparsity condition slightly, this rate can be proved to be much faster than the prediction error with the assumption $s\asymp r^{1+\delta}\phi_{r}\log r$. Hence, the bias term vanishes faster than the term $\sqrt{n/r^{2}\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x})}$ by the lower bound of $\gamma$, which completes the proof of Proposition 3.6.
@@
By Assumption 3.4, recall $\gamma>\frac{1}{1+\min_{j=0,...,q}(2m_{j}^{*})/(2m_{j}^{*}+t_{j})}$. For the sufficiently small $\varepsilon>0$, it is easy to see that
@@
Let $W_{b_{j}}$ be the number of subsamples $\mathcal{I}^{b_{j}}$ that contain index $i$, and define $g_{1,r}(\mathbf{z};\mathbf{x}_{*})=\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z},\mathbf{z}_{2},...,\mathbf{z}_{r})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})$. We have $W_{b_{j}}$ is independent of $\mathbf{z}_{i}$ and $\mathbb{E}W_{b_{j}}=B{{n-1}\choose{r-1}}/{{n}\choose{r}}$. Hence, the equation above involving the term $\mathbf{z}_{i}$ can be further expressed as
@@
$\frac{\sqrt{n}}{r\cdot\sqrt{\operatorname{{\rm Var}}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))}}\cdot\mathring{f}^{B}(\mathbf{x}_{*})~{\stackrel{{\scriptstyle d}}{{\rightarrow}}}~\mathcal{N}(0,1).$
By the definition of $\xi_{1,r}(\mathbf{x}_{*})$, it is easy to see that $\operatorname{{\rm Var}}(g_{1,r}(\mathbf{z}_{i};\mathbf{x}_{*}))=\xi_{1,r}(\mathbf{x}_{*})$, hence we have
@@
$\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big)$ $=\frac{1}{{{n}\choose{r}}}\Bigg({{n-1}\choose{r-1}}\sum_{i=1}^{n}T_{1}(\mathbf{z}_{i})+{{n-2}\choose{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})$
@@
$=\frac{1}{{{n}\choose{r}}}\Bigg({{n-2}\choose{r-2}}\sum_{i<j}T_{2}(\mathbf{z}_{i},\mathbf{z}_{j})+...+\sum_{i_{1}<...<i_{r}}T_{r}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{r}})\Bigg).$
Denote by $V_{k}=\operatorname{{\rm Var}}(T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}))$, we then calculate the value of $\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$, which we will show later this term is negligible. To calculate this term, we apply the equality $\mathbb{E}X_{n}^{2}=\mathbb{E}\operatorname{{\rm Var}}(X_{n}|Y_{n})+\mathbb{E}(\mathbb{E}X_{n}|Y_{n})^{2}$. Then the term $\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ can be written as
$\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ $=\underbrace{\mathbb{E}\big(\operatorname{{\rm Var}}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})|\mathbf{z}_{1},...,\mathbf{z}_{n}\big)\big)}_{J_{1}}$
@@
$J_{2}$ $=\frac{1}{{{n}\choose{r}}^{2}}\cdot\sum_{k=2}^{r}{{n-k}\choose{r-k}}^{2}\cdot{{n}\choose{k}}V_{k}$
$=\sum_{k=2}^{r}\frac{(n-k)!r!}{n!(r-k)!}{{r}\choose{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\cdot\sum_{k=2}^{r}{{r}\choose{k}}V_{k}$
$\leq\frac{r(r-1)}{n(n-1)}\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})).$
@@
$J_{1}$ $=\frac{1}{B^{2}}\sum_{k=2}^{r}\operatorname{{\rm Var}}(W_{i_{1}...i_{k}}){{n}\choose{k}}V_{k}$
$\leq\frac{1}{B}\sum_{k=2}^{r}\frac{{{n-k}\choose{r-k}}}{{{n}\choose{r}}}{{n}\choose{k}}V_{k}$
$=\frac{1}{B}\sum_{k=2}^{r}{{r}\choose{k}}V_{k}$
$\leq\frac{\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))}{B}.$
Here, the first inequality is by $\operatorname{{\rm Var}}(W_{i_{1}...i_{k}})\leq B\cdot\frac{{{n-k}\choose{r-k}}}{{{n}\choose{r}}}$ and the uncorrelation on different terms $T_{k}$, and the last inequality is by $\sum_{k=2}^{r}{{r}\choose{k}}V_{k}\leq\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$. We hence conclude that
@@
$\leq\bigg(\frac{r(r-1)}{n(n-1)}+\frac{1}{B}\bigg)\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r})),$
@@
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\mathbb{E}\big(\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-\mathring{f}^{B}(\mathbf{x}_{*})\big)^{2}$ $\leq\bigg(\frac{1}{n\xi_{1,r}(\mathbf{x}_{*})}+\frac{n}{Br^{2}\xi_{1,r}(\mathbf{x}_{*})}\bigg)\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},...,\mathbf{z}_{r}))$
@@
Here, we utilize the fact that $\widehat{f}^{b}(\mathbf{x}_{*})$ is bounded. The bound of $\widehat{f}^{b}(\mathbf{x}_{*})$ indicates that $\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*};\mathbf{z}_{1},\ldots,\mathbf{z}_{r}))$ is bounded. Additionally, we rely on Assumption 3.4 that $\liminf_{n\to+\infty}n\xi_{1,r}(\mathbf{x}_{*})\geq n^{1-\varepsilon-\gamma}\to+\infty$ and $\liminf_{n\to+\infty}Br^{2}\xi_{1,r}(\mathbf{x}_{*})/n\to+\infty$. From (C.6), we have
@@
Without loss of generality, we calculate the term $\widehat{V}_{1}$, and the terms $\widehat{V}_{i}$ with $i\neq 1$ are exactly the same. Consider $\widehat{V}_{1}^{2}$ and $B/n\to c$ for some constant $c>0$. Then by the inspiration of the proof of Lemma 13 of Wager and Athey (2018), we can define
@@
$\quad+\frac{r}{n}\sum_{k=1}^{r-1}\bigg(\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k+1}}}{{{n}\choose{k+1}}}\bigg)\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k+1}(\mathbf{z}_{1},\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}})$
$\quad+\frac{r}{n}\sum_{k=2}^{r}\bigg(\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k}}}{{{n}\choose{k}}}\bigg)\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}),$
@@
Further denote by $\widehat{\operatorname{{\rm Var}}}_{i}=\frac{1}{B-1}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}$, we have
$\widehat{R}_{B}=\frac{n(n-1)}{(n-r)^{2}}\frac{1}{B}\sum_{i=1}^{n}\widehat{\operatorname{{\rm Var}}}_{i}.$
@@
$R_{B}-\widehat{R}_{B}=\frac{n(n-1)}{(n-r)^{2}}\bigg(\underbrace{\sum_{i=1}^{n}2M_{i}\overline{\varepsilon}_{i}}_{\Delta_{1}}+\underbrace{\sum_{i=1}^{n}\overline{\varepsilon}_{i}^{2}-B^{-1}\operatorname{{\rm Var}}_{i}}_{\Delta_{2}}-\underbrace{\frac{1}{B}\sum_{i=1}^{n}\widehat{\operatorname{{\rm Var}}}_{i}-\operatorname{{\rm Var}}_{i}}_{\Delta_{3}}\bigg).$
Here, $\operatorname{{\rm Var}}_{i}=\operatorname{{\rm Var}}_{b}(Z_{b_{j}i})$ with variance condition on the dataset $\mathcal{D}_{n}$. We bound the three terms one by one. For the term $\Delta_{1}$, by the definition of $M_{i}$ and $\overline{\varepsilon_{i}}$, we have
@@
Conditional on the dataset $\mathcal{D}_{n}$, $\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}$ is independent with different $b_{j}$. It is easy to see that $\mathbb{E}_{b}\Delta_{1}=0$, and we give the calculation of $\operatorname{{\rm Var}}_{b}\Delta_{1}$. From the expression of $\Delta_{1}$, conditional on data set $\mathcal{D}_{n}$, $\varepsilon_{b_{j}i}$ are independent for different $b_{j}$, it is easy to see that
$\operatorname{{\rm Var}}_{b}\Delta_{1}\asymp\frac{\operatorname{{\rm Var}}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}}{B}.$
@@
$\operatorname{{\rm Var}}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}=\sum_{i=1}^{n}M_{i}^{2}\operatorname{{\rm Var}}_{b}\varepsilon_{b_{j}i}+\sum_{i\neq i^{\prime}}M_{i}M_{j}\operatorname{\rm Cov}_{b}(\varepsilon_{b_{j}i},\varepsilon_{b_{j}i^{\prime}}).$
From the definition of $M_{i}$, we can easily bound $M_{i}=O(r/n)$, for the covariance of $\varepsilon_{b_{j}i}$ and $\varepsilon_{b_{j}i^{\prime}}$, simple calculation of the subsampling randomness gives us that $\operatorname{{\rm Var}}_{b}\varepsilon_{b_{j}i}=O(\frac{r}{n}(1-\frac{r}{n})=O(r/n)$, and $\operatorname{\rm Cov}_{b}(\varepsilon_{b_{j}i},\varepsilon_{b_{j}i^{\prime}})=O(|\frac{r(r-1)}{n(n-1)}-\frac{r^{2}}{n^{2}})|=O(\frac{r}{n^{2}})$, we hence have
$\operatorname{{\rm Var}}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}$ $=O\bigg(n\cdot\frac{r^{2}}{n^{2}}\cdot\frac{r}{n}+n^{2}\cdot\frac{r^{2}}{n^{2}}\cdot\frac{r}{n^{2}}\bigg)$
@@
This indicates that $\operatorname{{\rm Var}}_{b}\Delta_{1}=O(\frac{r^{3}}{Bn^{2}})$. Combining the fact that $\mathbb{E}_{b}\Delta_{1}=0$, we conclude that $|\Delta_{1}|=O_{p}(\frac{r^{3/2}}{n\sqrt{B}})$.